OpenStudy (anonymous):

Equation pls 8.2p-33.4=1.7-3.5p

OpenStudy (anonymous):

Help pls

OpenStudy (debbieg):

Collect like terms to get all the p terms on one side (add 3.5p to both sides) and all the constant terms on the other (add 33.4 to both sides). Then you'll have {something}*p={something}, just divide both sides by the coefficient of p.

OpenStudy (mertsj):

1. Add 3.5p to both sides. 2. Add 33.4 to both sides 3. Divide both sides by the coefficient of p 4. You are done.

OpenStudy (anonymous):

Thanks y'all

OpenStudy (anonymous):

Can you try to draw it still not sure with my ans. So i can nail the next question

OpenStudy (mertsj):

What did you get for your answer?

OpenStudy (anonymous):

3p

OpenStudy (mertsj):

Drop the p

OpenStudy (mertsj):

p=3

OpenStudy (anonymous):

Did I get it right?

OpenStudy (mertsj):

Did you read what I posted?

OpenStudy (mertsj):

The answer is not 3p The answer is p=3

OpenStudy (anonymous):

Ohh okay thank you so much appreciate it

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Can you give me an idea for this one?

OpenStudy (anonymous):

x+y-a=b for y ___ z

OpenStudy (mertsj):

\[\frac{x+y}{z}-a =b\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

Yeah how'd you do that?

OpenStudy (mertsj):

With the equation button below this box.

OpenStudy (mertsj):

1. Add a to both sides 2. Multiply both sides by z 3. Subtract x from both sides.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

How will i do that with only letters not really a math kid :(

OpenStudy (mertsj):

Surely you can write +a _+a under both sides of the equation.

OpenStudy (anonymous):

I really don't know this one :(

OpenStudy (mertsj):